OpenStudy (anonymous):

3xsquared - 5x=0 how do I resoolve the x and the x squared

OpenStudy (candyme):

\[3x^2-5x=0\]Factor out an x \[x(3x-5)=0\] Set them equal to zero x=0 3x-5=0 3x=5 x=5/3 Hope this helps you!

OpenStudy (candyme):

So x=0, 5/3